the width of a rectangle is 4 inches less than than its length. The perimeter is 112 inches. Find the length and the width of the rectangle.
Answer by mangeshpjoshi(15)

formula for perimeter of rectangle
= 2(length + breadth)
let L= length so as per given instruction Breadth =B =L-4
so P= 112=2(L+L-4)
112/2= 2L-4
56=2L-4
60= 2L
L= 30
B= 30-4=26
thanks
faithfully
mangesh joshi